# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 8 Empirical Application
We adopt various ordered response models introduced in Section 3 to analyze the income category. The first model is the standard static model without any endogeneity. The second is the static model, while treating residence as an endogenous covariate. Residence is potentially endogenous since the choice of living area is typically endogenously determined and may be correlated with individuals’ unobserved ability or preference. The last model considers the dynamic model with one lagged dependent variable, allowing people’s income in current periods to depend on their income in the last period. All three models allow for individual fixed effects and do not impose any parametric distributions on time-changing shocks. Proposition 4 characterizes the identified set of the model coefficients for these three models using conditional moment inequalities. Similar to Section 7, we exploit the kernel-based CLR inference method to construct confidence intervals. The coefficient of the variable ‘residence’ is normalized to one. Table 6 reports the confidence intervals for the coefficients of the covariate ‘tenure’ and the lagged dependent variable (when applicable).
@@ A.1 Proof of Lemma 1
The identifying condition (7) for the parameter $\theta$ in Proposition 1 is given as
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}).$ (13)
Now we show that when $X_{it}\in\{a_{1},...,a_{K}\}$, the above condition is informative only at the $2K$ points $a\in\mathcal{Q}_{1}=\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma,z_{s}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{s}^{\prime}\beta+a_{K}^{\prime}\gamma\}$. Let $q_{k}\in\mathcal{Q}_{1}$ denote an element in $\mathcal{Q}$. Without loss of generalization, we can rank $q_{k}$ as $q_{1}\leq q_{2}\leq...\leq q_{2K}$.
When $a>q_{2K}$, then $a>z_{s}^{\prime}\beta+\max\limits_{k}{a_{k}^{\prime}\gamma}$. Therefore, the upper bound becomes $11-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})=$ 1 and the identifying condition holds for any $\theta$. Therefore, there is no information for $\theta$ when $a>q_{2K}$.
When $a<q_{1}$, similarly the lower bound is zero $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ and the inequality has no identifying power for $\theta$.
When $a\in(q_{k},q_{k+1})$ for some $k$, the identifying restriction at $a$ can be implied by the restriction at $a=q_{k+1}$. It is because the upper bound at $q_{k+1}$ is the same as $a$ since there does not exist any point $z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ that lies in $(q_{k},q_{k+1})$. The lower bound $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})$ is weakly larger at $q_{k+1}$ than $a$ since it is weakly increasing. Then, any parameter that satisfies condition (13) evaluated at $q_{k+1}$ also satisfies the one at $a$, so condition (13) evaluated at $a$ does not contain additional information for $\theta$.
@@ A.2 Proof of Theorem 1
To prove sharpness of the identified set $\Theta_{I,1}$, we seek to show that, for any $\theta\in\Theta_{I,1}$, there exists an underlying DGP characterized by the conditional distribution of $(v_{is},v_{it})\mid W_{ist}$) that produces the same observed choice probability and satisfies Assumption 1. The following construction is for one candidate parameter $\theta\in\Theta_{I,1}$. We set the fixed effect to be zero $\alpha_{i}=0$, so that $v_{it}:=-(\epsilon_{it}+\alpha_{i})=-\epsilon_{it}$.
Let $P_{st}\left(j,k\mid w_{st}\right):=\Pr\left(Y_{is}=j,Y_{it}=k\mid W_{ist}=w_{st}\right)$ denote the observed joint probability of selecting $j$ at period $s$ and selecting $k$ at period $t$. The first requirement for sharpness is that the constructed distribution of $(v_{is},v_{it})\mid W_{ist}$ matches the observed joint choice probability according to the model: given $W_{ist}=w_{st}$,
@@
$P_{st}\left(1,1\mid w_{st}\right)$ $=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$ (14)
@@
$P_{st}\left(1,0\mid w_{st}\right)$ $=\Pr(v_{is}\leq w_{s}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
$P_{st}\left(0,1\mid w_{st}\right)$ $=\Pr(v_{it}\leq w_{t}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
and $P_{st}\left(0,0\mid w_{st}\right)$ can be matched automatically when the above conditions hold.
Another condition for sharpness requires the marginal distribution of $v_{it}\mid W_{ist}$ to satisfy the following stationarity condition in Assumption 1:
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})=F_{v_{s}\mid Z_{st}}(a\mid z_{st})\quad\forall a\in\mathcal{R}.$ (15)
We can focus on a reduced problem, which is to find a marginal distribution $v_{it}\mid W_{ist}$ that matches the marginal choice probability and satisfies the stationarity assumption. Let $P_{t}(w_{st}):=\Pr(Y_{it}=1\mid W_{ist}=w_{st})$ and $P_{s}(w_{st}):=\Pr(Y_{is}=1\mid W_{ist}=w_{st})$ denote the observed marginal choice probability at $t$ and $s$, given as follows:
$P_{t}(w_{st})=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta\mid W_{ist}=w_{st}\right),$ (16)
$P_{s}(w_{st})=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta\mid W_{ist}=w_{st}\right).$
If we can construct a marginal distribution satisfying conditions (15) and (16), then the joint choice probability in (14) is matched by setting the joint distribution of $(v_{is},v_{it})\mid W_{ist}$ at $(w_{s}^{\prime}\theta,w_{t}^{\prime}\theta)$ as $P_{st}\left(1,1\mid w_{st}\right)=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right)$. This construction is feasible since there is no other assumptions on the joint distribution $(v_{is},v_{it})\mid W_{ist}$. Moreover, other choice probabilities such as $P_{st}\left(1,0\mid w_{st}\right)$ can be computed using the marginal probability $P_{s}(w_{st})$ subtracting $P_{st}\left(1,1\mid w_{st}\right)$.
Now, for any $\theta\in\Theta_{I,1}$, our objective is to construct the distribution of $v_{it}\mid W_{st}$ and $v_{is}\mid W_{st}$ such that conditions (15) and (16) hold. Consider that the endogenous covariate $X_{it}$ only takes $K$ values: $X_{it}\in\mathcal{X}=\{a_{1},...,a_{K}\}$ for any $t\leq T$.  [^10] The covariate $X_{it}$ is allowed to have different support across time $t$, then $\mathcal{X}$ is the union of the support of $X_{it}$ over all periods. Given $Z_{it}=z_{t}$, the covariate index $z_{t}^{\prime}\beta+x_{t}^{\prime}\gamma$ belongs to the $K$ points $\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma\}$. Without loss of generalization, we can rank the $K$ points as $\alpha_{1}^{\prime}\gamma\leq a_{2}^{\prime}\gamma\leq...\leq a_{K}^{\prime}\gamma$ since we can reindex the $K$ values of $\{a_{1},...,a_{K}\}$ such that this ranking is satisfied. Let $m_{k}:=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$ denote the $K$ points for which $w_{t}^{\prime}\theta$ can take, and $n_{k}:=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ for $1\leq k\leq K$ denote the $K$ points $w_{s}^{\prime}\theta$ can take. Let $\bar{m}=\max\{z_{t}^{\prime}\beta,z_{s}^{\prime}\beta\}+a_{K}^{\prime}\gamma+\epsilon$ for some $\epsilon>0$.
We first look at the stationarity condition in (15). Section 2.2 derives bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, given as follows:
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\leq\sum_{x_{st}\in\mathcal{X}}\Big{\{}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\geq a\}+\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta<a\}\Big{\}}\Pr(X_{st}=x_{st}\mid z_{st})$
$:=U_{t}(a\mid z_{st}),$
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\geq\sum_{x_{st}\in\mathcal{X}}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\leq a\}\Pr(X_{st}=x_{st}\mid z_{st}):=L_{t}(a\mid z_{st}).$
Given the definition of the identified set $\Theta_{I,1}$, for $\theta\in\Theta_{I,1}$, the following condition holds for any $a\in\mathcal{R}$,
$\max_{s,t}\{L_{s}(a\mid z_{st}),L_{t}(a\mid z_{st})\}\leq\min_{s,t}\{U_{t}(a\mid z_{st}),U_{s}(a\mid z_{st})\},$
Let $\tilde{L}(a\mid z_{st})=\max_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}$, the stationarity condition (15) is satisfied by constructing the conditional distribution of $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ as follows:
@@
$F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})=\left\{\begin{aligned} &0\ &\text{if }a<m_{1},\\ \tilde{L}&(a\mid z_{st})\ &\text{if }m_{1}\leq a<\bar{m},\\ &1\ &\text{if }\bar{m}\leq a.\end{aligned}\right.$ (17)
The above construction guarantees that the marginal distribution $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})$ lies between the bounds $[L_{s}(a\mid z_{st}),U_{s}(a\mid z_{st})]$. From the definition of $L_{t}(a)$, we know that $L_{t}(a\mid z_{st})\geq L_{t}(a^{\prime}\mid z_{st})$ for any $a\geq a^{\prime}$. Therefore, the monotonicity of the above distribution is satisfied as $\max\limits_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}\geq\max\limits_{s,t}\{L_{t}(a^{\prime}\mid z_{st}),L_{s}(a^{\prime}\mid z_{st})\}$ for any $a\geq a^{\prime}$. Additionally, the above construction satisfies the stationarity assumption given $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ for any $a$.
We transform the two requirements of sharpness into finding a distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ to generate the marginal distribution $F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ in (17) and also match the marginal choice probability given in condition (16).
Construction
We focus on the construction for $v_{it}\mid W_{ist}$, and the construction for $v_{is}\mid W_{ist}$ at period $s$ is similar so it is omitted here. The construction proceeds in four steps. We first constructs a distribution to match the lower and upper bounds $L_{t}(a\mid z_{st})$ and $U_{t}(a\mid z_{st})$. Then, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $\{m_{1},...,m_{K}\}$ with $m_{k}=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$. The third step look at the construction at the points $\{n_{1},...,n_{K}\}$ with $n_{k}=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$. The final step examines the points $a$ that are not in $\{m_{1},...,m_{K},n_{1},...,n_{K}\}$.
@@
We first show that the bounds $U_{t}(a\mid z_{st})$ and $L_{t}(a\mid z_{st})$ can be achieved under a distribution of $v_{it}\mid W_{ist}$ that matches the marginal choice in (16). From its definition, the upper bound $U_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})$:
$F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<m_{1},\\ &P_{t}(w_{st})\qquad&\text{if }m_{1}\leq a<w_{t}^{\prime}\theta+\epsilon,\\ &1\qquad&\text{if }w_{t}^{\prime}\theta+\epsilon\leq a,\end{aligned}\right.$
where $\epsilon>0$ is used to satisfy the right-continuity property of the distribution.
Similarly, the lower bound function $L_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$:

### v3

@@ 5 Empirical Application
We adopt various ordered response models introduced in Section 3.2 to analyze the income category. The first model is the standard static model without any endogeneity. The second is the static model, while treating residence as an endogenous covariate. Residence is potentially endogenous since the choice of living area is typically endogenously determined and may be correlated with individuals’ unobserved ability or preference. The last model considers the dynamic model with one lagged dependent variable, allowing people’s income in current periods to depend on their income in the last period. All three models allow for individual fixed effects and do not impose any parametric distributions on time-changing shocks. Proposition 7 characterizes the identified set of the model coefficients for these three models using conditional moment inequalities. Similar to Section 4, we exploit the kernel-based CLR inference method to construct confidence intervals. The coefficient of the variable ‘residence’ is normalized to one. Table 6 reports the confidence intervals for the coefficients of the covariate ‘tenure’ and the lagged dependent variable (when applicable).
@@ A.1 Proof of Proposition 2
Clearly, $\Theta_{I}\subseteq\Theta^{disc}$. Below we show $\Theta^{disc}\subseteq\Theta_{I}$ when $X_{it}$ is discrete. Suppose that $\theta$ satisfies condition (10) at all
$c\in{\cal C}\left(\theta\right):=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\}$
for any realization $z=\left(z_{1},...,z_{T}\right).$ We seek to show that $\theta$ must also satisfy condition (10) for any $c\in\mathcal{R}\backslash{\cal C}\left(\theta\right)$. Without loss of generality, we order elements in ${\cal C}\left(\theta\right)$ from the smallest to the largest as
$\overline{c}_{1}\leq\overline{c}_{2}\leq...\leq\overline{c}_{KT}.$
For $c<\overline{c}_{1}$, we must have
$\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)\equiv 0,$
so (10) holds trivially. Similarly, for $c>\overline{c}_{KT}$, we must have
$\mathbb{P}\left(\left.Y_{is}=1,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\equiv 0,$
so (10) again holds trivially. For any $c$ s.t. $\overline{c}_{j}<c<\overline{c}_{j+1}$ for some $j$, we have
$z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\quad\Leftrightarrow\quad z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}$
and
$z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\quad\Leftrightarrow\quad z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}.$
which implies
$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)=$ $\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}\right|Z_{i}=z\right)$ (26)
and
$\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$ $=\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}\right|Z_{i}=z\right)$
$\leq\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right),$
or equivalently,
$1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}\right|Z_{i}=z\right).$ (27)
Since (10) holds at $\overline{c}_{j}$, we have
$\max_{t}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}\right|Z_{i}=z\right)\leq 1-\max_{s}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right).$
Combining the above with (26) and (27), we have
$\max_{t}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\leq 1-\max_{s}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right).$
@@ A.2 Proof of Theorem 1
For shorter notation we write $W_{it}:=\left(Z_{it},X_{it}\right)$ for the combination of the exogenous covariates $Z_{it}$ and endogenous covariates $X_{it}$. Correspondingly, we write $W_{i}\equiv\left(W_{i1},...,W_{iT}\right)$, $W_{i}^{{}^{\prime}}\theta_{0}\equiv Z_{i}^{{}^{\prime}}\beta_{0}+X_{i}^{{}^{\prime}}\gamma_{0}$ and the lower cases $w\equiv\left(w_{1},...,w_{T}\right)$ for realizations.
We first clarify the rigorous meaning of “sharpness” in Theorem 1 through the following definition.
Definition 1.
We say that $\Theta_{I}$ is sharp under model (5) and Assumption 1 if, for any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exist well-defined latent random variables $\left(\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ such that:
  • Assumption 1 (partial stationarity) is satisfied, i.e.,
    $\left.\epsilon_{it}^{*}\sim\epsilon_{is}^{*}\right|Z_{i},\alpha_{i}^{*},\ \forall t,s=1,...,T.$
  • (CCP-J) $\left(\theta,\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ are observationally equivalent to $\left(\theta_{0},\epsilon_{i},\alpha_{i}\right)$, i.e., formally, $\left(\theta,\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ produces the following conditional choice probabilities under model (5):
    $\mathbb{P}\left(v_{it}^{*}\leq w_{t}^{{}^{\prime}}\theta\ \forall t\text{ s.t. }y_{t}=1,\ v_{is}^{*}>w_{s}^{{}^{\prime}}\theta\ \forall s\text{ s.t. }y_{s}=0\mid w\right)=p\left(y\mid w\right),$ (28)
    where $v_{it}^{*}:=-\left(\epsilon_{it}^{*}+\alpha_{i}^{*}\right)$ and $p\left(\left.\cdot\right|w\right)$ denotes the true conditional probability
    $p\left(\left.y\right|w\right)$ $:=\mathbb{P}\left(Y_{it}=y_{t}\,\ \forall t=1,...,T\mid W_{i}=w\right)$
@@
    $\equiv\mathbb{P}\left(v_{it}\leq w_{t}^{{}^{\prime}}\theta_{0}\ \forall t\text{ s.t. }y_{t}=1,\ v_{is}>w_{s}^{{}^{\prime}}\theta\ \forall s\text{ s.t. }y_{s}=0\mid W_{i}=w\right),$
    for any outcome realization $y\equiv\left(y_{1},...,y_{T}\right)\in\left\{0,1\right\}^{T}$, given any realization $W_{i}=w$.
We prove Theorem 1, i.e., the sharpness of the identified set $\Theta_{I}$ under discreteness of $X_{it}$ by, for any candidate parameter $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, we construct the distribution of $\left(\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ to satisfy the above requirements.
Proof.
Set $\alpha_{i}^{*}\equiv 0$ and $\epsilon_{i}^{*}:=-v_{i}^{*}$. Then the conclusion follows from Lemma 1 and 2 below. ∎
Lemma 1 (Per-Period Construction).
When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for all $t=1,...,T$, then for any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exist well-defined latent random variables $v_{i1}^{*},...,v_{iT}^{*}$ with marginal CDFs $F_{1}^{*},...,F_{T}^{*}$ such that
$F_{t}^{*}\left(\left.\cdot\right|Z_{i}=z\right)=F_{s}^{*}\left(\left.\cdot\right|Z_{i}=z\right)$ (29)
and
$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)=p_{t}\left(w\right),\quad\forall t,\forall w,$ (30)
where
$p_{t}\left(w\right):=\mathbb{P}\left(\left.Y_{it}=1\right|W_{i}=w\right).$
Proof.
For any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, below we show how to construct $v_{i1}^{*},...,v_{iT}^{*}$, or equivalently, the conditional CDFs $F_{1}^{*}\left(\left.c\right|W_{i}=w\right),...,F_{T}^{*}\left(\left.c\right|W_{i}=w\right)$ for each realization $w$ and each $c\in\mathcal{R}$ so that (i) condition (29) is satisfied so that partial stationarity holds; and (ii) condition (30) is satisfied so that per-period marginal CCPs are matched.
Fix a specific realization of the exogenous covariates at $z\equiv\left(z_{1},...,z_{T}\right)$. We construct the (conditional) CDF $F_{t}^{*}$ of $v_{it}^{*}$ for each $t=1,...,T$ and each given $z$ in the following manner.
Define
$L_{t}\left(\left.c\right|z\right)$ $:=\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right),$
$U_{t}\left(\left.c\right|z\right)$ $:=1-\mathbb{P}\left(\left.Y_{it}=0,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right),$
and
$\overline{L}\left(\left.c\right|z\right):=\max_{s}L_{s}\left(\left.c\right|z\right),\quad$ $\underline{U}\left(\left.c\right|z\right):=\min_{s}U_{s}\left(\left.c\right|z\right).$
@@
Since $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, by 10 we have,
$\overline{L}\left(\left.c\right|z\right)\leq\underline{U}\left(\left.c\right|z\right),\quad\forall c\in\mathcal{R}.$
Observe that both $\overline{L}\left(\left.c\right|z\right)$ and $\underline{U}\left(\left.c\right|z\right)$ are weakly increasing in $c$.
[…]
${\cal C}:=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:t=1,...,T,k=1,...,K\right\}.$
Let $\delta>0$ be a sufficiently small constant [^10] The small positive constant $\delta>0$ is used to ensure the right continuity of CDFs defined afterwards. Let $\underline{\delta}$ to be smallest distance between two distinct points in ${\cal C}$. If $\underline{\delta}>0$, then we may set $\delta:=\underline{\delta}/2$. If $\underline{\delta}=0$, then $\delta$ can be set as any positive number, say, $\delta:=1$., and define
$\underline{c}:=\min{\cal C},\quad\overline{c}:=\max{\cal C}+\delta.$
The parametrix index $w_{t}^{{}^{\prime}}\theta$ must lie within the interval $\left[\underline{c},\overline{c}\right)$ across all $t$ for all possible realization of $x$.
For each $t=1,...,T,$ we show how to construct $v_{t}^{*}$ with CDF $F_{t}^{*}$ so that
@@
$F_{t}^{*}\left(\left.c\right|z\right)\equiv\begin{cases}0,&\text{if }c<\underline{c},\\ \overline{L}\left(\left.c\right|z\right),&\text{if }\underline{c}\leq c<\overline{c},\\ 1,&\text{if }c\geq\overline{c},\end{cases}$ (31)
and
$F_{t}^{*}\left(\left.w_{t}^{\prime}\theta\right|w\right)=p_{t}\left(w\right).$ (32)
Clearly, partial stationarity (29) will be satisfied under (31), the right-hand side of which does not depend on the time index $t$. Furthermore, (32) is the same as (30), i.e., the marginal CCPs will be matched for each $t$. ∎
Proof.
@@
We construct the conditional CDF of $v^{*}|W_{i}=w$ using two auxillary CDFs $F_{t}^{L}$ and $F_{t}^{U}$, defined by
$F_{t}^{L}\left(\left.c\right|w\right)$ =$=\begin{cases}0,&c<w_{t}^{{}^{\prime}}\theta,\\ p_{t}\left(w\right),&w_{t}^{{}^{\prime}}\theta\leq c<\overline{c}_{t},\\ 1,&c\geq\overline{c}_{t},\end{cases}$
and